OpenStudy (anonymous):

what is the slope of the line whose equation is 2y=5x+4 ?

OpenStudy (mitu12):

5/2 y = mx + b, so just divide everything by 2 to get y = (5/2)x + 2, so you see that m = 5/2.

OpenStudy (anonymous):

Y=5/2x+2

OpenStudy (anonymous):

How would you solve it step by step @mitu12

OpenStudy (mitu12):

dividing everything by 2 will get you y=(5\2)x+2

OpenStudy (mitu12):

then y=mx+b your x will be the slope

OpenStudy (anonymous):

so i what would i divide first ?

OpenStudy (mitu12):

the 5 which will give you (5\2) leave it as an improper fraction

OpenStudy (mitu12):

after that if you divide 4\2 that will give you 2

OpenStudy (anonymous):

i dont get it, would you please solve

OpenStudy (mitu12):

\[2y=5x+4 \] \[5\2x+4/2\]

OpenStudy (anonymous):

equals to 2

OpenStudy (mitu12):

4 divided by 2 =2

OpenStudy (mitu12):

you leave 5\2 as an improper fraction because 2 can not be evenly divided by 5

OpenStudy (mitu12):

5/2 will be you x and 2 will be your b according to y=mx+b

OpenStudy (anonymous):

and what is y= ?

OpenStudy (mitu12):

y= (5\2)x +2